OpenStudy (anonymous):

tan x = .39 and cos x = .93, what is sin x?

OpenStudy (anonymous):

tan x = sin x / cos x

OpenStudy (anonymous):

0.39 = sin x/ 0.93 sin x = 0.39 * 0.93

OpenStudy (anonymous):

great thanks so much!

OpenStudy (anonymous):

tan x = (sin x)/(cos x) 0.39 = (sin x)/0.93 *0.93 *0.93 0.3627= sin x